# reference.bib

### v1

@@ 6. Relation to the class number
We now assert we can take $\xi_{i}\in\mathcal{P}_{S}=\{y=(y_{1},y,y_{2})\in\mathbb{R}^{n+2}\mid y_{1}>0,\phi_{0}[y]>0\}$ of Section 2 for all $i$. The second condition is clear, as $S_{0}[\xi_{i}]=-2D/q>0$ because we take $D<0$. For the first one, we can always mutliply with $-1_{n+2}\in\Gamma(L_{0})$ and the assertion follows.
Let now $F\in S_{k}(\Gamma_{S})$. For an element $\gamma^{\prime}\in\Gamma_{S}$ of the form $\gamma^{\prime}=\begin{pmatrix}1&0&0\\ 0&\gamma&0\\ 0&0&1\end{pmatrix}$ we have $F(\gamma^{\prime}\langle Z\rangle)=F(\gamma Z)$.
In particular this then implies that $A(\gamma r)=A(r)$ for all $r\in L_{0}^{*}$ and $\gamma\in\Gamma^{+}(L_{0})$, where we use the Fourier expansion of $F$ of Definition 2.5:
$F(Z)=\sum_{r\in L_{0}^{*}\cap\mathcal{P_{S}}}A(r)e(r^{t}S_{0}Z).$
Let us now consider the subgroup of elements $\gamma\in\Gamma(L_{0})$ such that $\begin{pmatrix}1&0&0\\ 0&\gamma&0\\ 0&0&1\end{pmatrix}\in\Gamma_{S}$. The only extra condition we need is that $\gamma\in G^{*,0}_{\mathbb{R}}$ (the connected component of the identity) also.
By [krieg_integral_orthogonal, page 180], we see that if $\gamma=\begin{pmatrix}*&*&*\\ x&*&y\end{pmatrix}$, the condition we need is $x+y>0$. Hence, for any $\gamma\in\Gamma(L_{0})$,
$\begin{pmatrix}1&0&0\\ 0&\gamma&0\\ 0&0&1\end{pmatrix}\in\Gamma_{S}\textup{ or }\begin{pmatrix}1&0&0\\ 0&-\gamma&0\\ 0&0&1\end{pmatrix}\in\Gamma_{S}.$
In the second case, we obtain for $\xi\in\Xi_{d}$ of Lemma 6.1, that:
$A\left(\xi\right)=A(\gamma\xi_{i})=A\left(-\xi_{i}\right)=0,$
because $A(r)=0$ unless $r\in L_{0}^{*}\cap\mathcal{P}_{S}$ (cf. Definition 2.5).
Define now the following quantities:
$n(\xi_{i};d):=\#\left\{s\in\mathbb{Z}^{n}/dS\mathbb{Z}^{n}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd}\textup{ and }\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\gamma\in\Gamma^{+}(L_{0})\right\},$
$m(\xi_{i};d):=\#\left\{s\in\mathbb{Z}^{n}/dS\mathbb{Z}^{n}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd}\textup{ and }\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\Gamma(L_{0})\backslash\Gamma^{+}(L_{0})\right\}.$
From the above considerations, Lemma 6.1 and Proposition 5.2, we may write, for $(N,D)=1$:
$\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle=\sum_{i=1}^{h}\sum_{0<d|N}d^{k-(n+1)}\left[n(\xi_{i};d)A\left(\frac{N}{d}\xi_{i}\right)+m(\xi_{i};d)A\left(-\frac{N}{d}\xi_{i}\right)\right]=$
$=\sum_{i=1}^{h}\sum_{0<d|N}d^{k-(n+1)}n(\xi_{i};d)A\left(\frac{N}{d}\xi_{i}\right).$
In particular, we arrive at the following proposition:
@@
Let $\mathcal{P}$ be a sufficiently large finite set of primes, which includes the prime factors of $D$. Let also
@@
Define also:
for $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. We then have that:
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=\sum_{i=1}^{h}\zeta_{\xi_{i}}(s-k+n+1)D_{F,\xi_{i},\mathcal{P}}(s),$
where $D_{F,\xi_{i},\mathcal{P}}(s)=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}A(N\xi_{i})N^{-s}$.
@@ 7. Relation to Sugano’s Theorem
In order to now obtain an Euler product, we will make use of the main theorem of Sugano in his paper [sugano]. We first need some setup:
@@
We remind ourselves here that $G_{\infty}^{0}$ acts transitively on $\mathcal{H}_{S}$ (cf. Section 2). Let $\mathcal{Z}_{0}$ denote any point of $\mathcal{H}_{S}$ with real part 00 0 and denote by $K_{\infty}$ its stabiliser in $G_{\infty}^{0}$. Then, we have that $G_{\infty}^{0}/K_{\infty}\cong\mathcal{H}_{S}$.
@@
$\bm{F}(g_{f};\mathcal{Z})=\bm{F}(g_{\infty}g_{f})j(g_{\infty},\mathcal{Z}_{0})^{k},$
@@
$\bm{F}(g_{f};\gamma\langle\mathcal{Z}\rangle)=j(\gamma,\mathcal{Z})^{k}\bm{F}(g_{f};\mathcal{Z}).$
for all $\gamma\in\Gamma(g_{f})$ and $\mathcal{Z}\in\mathcal{H}_{S}$. Now, if $X\in V_{0}$, define the element $\gamma_{X}\in G$ by:
$\gamma_{X}:=\begin{pmatrix}1&-X^{t}S_{0}&-\frac{1}{2}S_{0}[X]\\ 0&1_{q}&X\\ 0&0&1\end{pmatrix}.$
Now, the holomorphic function $\bm{F}(g_{f};\mathcal{Z})$ is invariant under $\mathcal{Z}\longmapsto\mathcal{Z}+X$ for $X$ in the lattice:
@@
Hence, every such function then admits a Fourier expansion of the form:
$\bm{F}(g_{f};\mathcal{Z})=\sum_{\begin{subarray}{c}r\in\hat{L}(g_{f})\\ ir\in\mathcal{H}_{S}\end{subarray}}a(g_{f};r)e[2\phi_{0}(r,Z)],$
@@
Finally, let us introduce adelic Fourier coefficients. Let $\chi=\prod_{p\leq\infty}\chi_{p}$ be a character of $\mathbb{Q}_{\mathbb{A}}$ such that $\chi_{\mid\mathbb{Q}}=1$ and $\chi_{\infty}(x)=e[x]$ for all $x\in\mathbb{R}$. For $\eta\in V_{0}$ and $g\in G_{\mathbb{A}}$, we define:
[…]
@@
$\bm{F}_{\chi}(g_{\infty}g_{f};\eta)=a(g_{f};\eta)j(g_{\infty};\mathcal{Z}_{0})^{-k}e[2\phi_{0}(\eta,g_{\infty}\langle\mathcal{Z}_{0}\rangle)].$
We then have from the above definitions (see [sugano, (1.14)]):
@@
From [shimura2004arithmetic, Lemma 17.6, (i)], there is a bijection between $\mathfrak{S}(K_{f})$ and $S_{k}(\Gamma_{S})$, which is given by $\bm{F}\longmapsto F(\mathcal{Z}):=\bm{F}(g_{\infty})j(g_{\infty},\mathcal{Z}_{0})^{k}$, where $g_{\infty}\in G_{\infty}^{0}$ is chosen so that $g_{\infty}\langle\mathcal{Z}_{0}\rangle=\mathcal{Z}$. This follows by the fact that the class number of $G_{\mathbb{Q}}$ is 11 11, which we can show in the same way as in Lemma 6.2.
In particular, it then follows that we can specify our $F\in S_{k}(\Gamma_{S})$ of previous sections with $\bm{F}(id,\mathcal{Z})$. In that case, i.e. when $g_{f}=(id,id,\cdots,id)$, which we will just denote by $id$, we have that $L(g_{f})=L=\mathbb{Z}^{n+2}$. Now, $a(id,r)=A(r)$ for all $r\in L_{0}^{*}$, so in particular:
@@
for all $i$, where $\xi_{i}$’s are the representatives for $L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})$.
Let now $\{u_{i}\}_{i=1}^{h}$ denote a complete system of representatives for $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ such that $u_{i,\infty}=1$ for all $i=1,\cdots,h$. We then prove the following lemma:
@@
Let $\xi$ be a specific element of $L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]$, as we have specified it right after Lemma 6.2. If now $\xi_{i}$’s and $u_{i}$’s are as above, we have
@@
By the definition of the $\xi_{i}$’s and Witt’s theorem ([shimura2004arithmetic, Lemma 1.5 (ii)]), there is $\alpha\in G_{\mathbb{Q}}^{*}$ such that $\xi_{i}=\alpha\xi$. By the correspondence given by Shimura in Proposition 6.2, we get $\alpha^{-1}L_{0}=u_{i}L_{0}$. This then implies $\alpha u_{i}L_{0}=L_{0}$, so $\alpha u_{i}\in C$ (we remind ourselves that $C=\{x\in G^{*}_{\mathbb{A}}\mid xL_{0}=L_{0}\}$). By definition, we then get $\alpha u_{i}\in K_{f}$. Hence, we obtain:
$\bm{F}_{\chi}(u_{i};\xi)=\bm{F}_{\chi}(u_{i};\alpha^{-1}\xi_{i})=\bm{F}_{\chi}\left(\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix}u_{i};\xi_{i}\right)=\bm{F}_{\chi}\left(\left(\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix},id\right);\xi_{i}\right),$
where $\left(\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix},id\right)$ denotes the element of $G_{\mathbb{A}}$ with infinity part $\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix}$. The second equality above follows from the second bullet point and the third equality from the first bullet point above. By the property of $\bm{F}_{\chi}$, just before the bullet points, we obtain:
$a(u_{i};\xi)e[2\phi_{0}(\xi,Z_{0})]=a(id,\xi_{i})j\left(\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix},Z_{0}\right)^{-k}e\left[2\phi_{0}\left(\xi_{i},\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix}\langle Z_{0}\rangle\right)\right].$
But $j\left(\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix},Z_{0}\right)=1$ and $\begin{pmatrix}1\\ &\alpha\\ &&1\end{pmatrix}\langle Z_{0}\rangle=\alpha Z_{0}$ which then gives that the right hand side above equals:
$a(id;\xi)e[2\phi_{0}(\xi_{i},\alpha Z_{0})]=a(id;\xi)e[2\phi_{0}(\alpha\xi,\alpha Z_{0})]=a(id;\xi)e[2\phi_{0}(\xi,Z_{0})],$
@@
For each prime $p$, let $M(g_{f};\xi)_{p}=H(\xi)_{p}\cap g_{f}K_{f}g_{f}^{-1}$ and $M(g_{f};\xi)_{f}=\prod_{p}M(g_{f};\xi)_{p}$. Let us now recall the main theorem in the paper of Sugano ([sugano, Theorem 1]):
@@
Let $g_{f}=(id,\cdots,id)\in G_{\mathbb{A},f}$ be fixed and $\xi\in\hat{L}(g_{f})$ with $i\xi\in\mathcal{H}_{S}$. We identify $F\in S_{k}(\Gamma_{S})$ with $\bm{F}(g_{f};Z)\in\mathfrak{S}_{k}(K_{f})$ as above. Denote by $V(g_{f};\xi)$ the space of functions on $H(\xi)_{\mathbb{A}}$, which are left $H(\xi)_{\mathbb{Q}}$ and right $H(\xi)_{\mathbb{A}}\cap C$ invariant. Let, as above, $\{u_{i}\}_{i=1}^{h}$ denote a complete system of representatives for $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ such that $u_{i,\infty}=1$ for all $i=1,\cdots,h$. Let $\mathcal{P}$ be a sufficiently large set of primes such that $\bm{F}$ is a simultaneous eigenfunction of the Hecke pair $\mathcal{H}_{p}=(G_{p},K_{p})$ for all $p\notin\mathcal{P}$ and $f\in V(g_{f};\xi)$ be a simultaneous eigenfunction of the Hecke algebra $\mathcal{H}_{p}^{\prime}:=(H(\xi)_{p},M(g_{f};\xi)_{p})$ for all $p\notin\mathcal{P}$. Then, we have:
@@
$\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}\mu(\xi)^{-1}\sum_{i=1}^{h}A(N\xi_{i})\frac{\overline{f}(u_{i})}{e(\xi)_{i}}N^{-(s+k-\frac{n+2}{2})}=\left(\mu(\xi)^{-1}\sum_{i=1}^{h}A(\xi_{i})\frac{\overline{f}(u_{i})}{e(\xi)_{i}}\right)L_{\mathcal{P}}(F;s)\times$
×$\times L_{\mathcal{P}}\left(\overline{f};s+\frac{1}{2}\right)^{-1}\times(\textup{zeta})^{-1},$
where $(\textup{zeta})=\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s)&\textup{if }n\textup{ even}\end{cases}$.
@@
The definitions of the Hecke algebras can be found in [sugano, Section 2]. $L(-,s)$ denotes the standard $L-$function attached to orthogonal modular forms, as this is defined in [sugano, Paragraph 4-1]. Also, $\zeta(s)$ denotes the Riemann zeta function and for any $L-$function, we write $L_{\mathcal{P}}$ for the Euler product not containing the primes in $\mathcal{P}$. Also, we define $e(\xi)_{i}=\#\{H(\xi)_{\mathbb{Q}}\cap M(u_{i}g_{f};\xi)_{f}\}$ for $1\leq i\leq h$ and $\mu(\xi)=\sum_{i=1}^{h}e(\xi)^{-1}_{i}$.
Remark 6.
We have adapted the above theorem in our setting, meaning we have taken $g_{f}=(id,id,\cdots,id)$ and then substituted $a(u_{i};N\xi)$ with $A(N\xi_{i})$ in the original form of [sugano, Theorem 1]. We also note that $H(\xi)_{\infty}M(g_{f};\xi)_{f}=H(\xi)_{\mathbb{A}}\cap C$.
By setting then $g(u_{i})=\frac{\overline{f}(u_{i})}{e(\xi)_{i}}\mu(\xi)^{-1}$ and $A_{g}=\sum_{i=1}^{h}g(u_{i})A(\xi_{i})$, the formula in Theorem 7.1 becomes:
@@
$(\textup{zeta})\times L_{\mathcal{P}}\left(\overline{f};s+\frac{1}{2}\right)\sum_{i=1}^{h}g(u_{i})D_{F,\xi_{i},\mathcal{P}}\left(s+k-\frac{n+2}{2}\right)=A_{g}L_{\mathcal{P}}(F;s),$
@@
where $D_{F,\xi_{i},\mathcal{P}}$ is the Dirichlet series appearing in Proposition 6.3. This is true for any simultaneous eigenfunction $f$ (hence $g$ as well) of the Hecke algebras $H_{p}^{\prime}=(H(\xi)_{p},M(g_{f},\xi)_{f})$ for all $p\notin\mathcal{P}$. Our aim is to try to invert it so that we solve for $D_{F,\xi_{i},\mathcal{P}}$.
Now, any such $f$ is a $\mathbb{C}-$valued function on $H(\xi)_{\mathbb{A}}$, which is left $H(\xi)_{\mathbb{Q}}$ invariant and right $H(\xi)_{\infty}M(g_{f};\xi)_{f}$ invariant. By what we are given, we have that
@@
where $D=H(\xi)_{\infty}M(g_{f};\xi)_{f}$. We note $D$ is an open subgroup of $H(\xi)_{\mathbb{A}}$ and $D\cap H(\xi)_{f}$ is compact. By [shimura2004arithmetic, Lemma 17.6], there is a correspondence

### v2

@@ 7. Relation to the class number
We now assert we can take
$\xi_{i}\in\mathcal{P}_{S}=\{y^{\prime}=(y_{1},y,y_{2})\in\mathbb{R}^{n+2}\mid y_{1}>0,\ \phi_{0}[y^{\prime}]>0\}$
of Section 2 for all $i$. The second condition is clear, as $S_{0}[\xi_{i}]=-2D/q>0$ because we take $D<0$. For the first one, we can always multiply with $\textup{diag}(-1,1_{n},-1)\in\Gamma(L_{0})$ and the assertion follows.
[…]
Therefore, from [21, p. 26], we have $A(\xi)=A(\gamma\xi_{i})=A(\xi_{i})$. Define now
$n(\xi_{i};d):=\#\left\{s\in\mathbb{Z}^{n}/dS\mathbb{Z}^{n}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\Gamma^{+}(L_{0})\right\}.$
From the above considerations, Lemma 7.1 and Proposition 6.1, we may write, for $(N,D)=1$
$\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle=\sum_{i=1}^{h}\sum_{0<d|N}d^{k-(n+1)}n(\xi_{i};d)A\left(\frac{N}{d}\xi_{i}\right).$
In particular, we arrive at the following Proposition.
@@
Let $(D,r)\in\widetilde{\textup{supp}}(L,\sigma)$. Let $\mathcal{P}$ be a finite set of primes, which includes the prime factors of $D$. Let
which converges absolutely for $\textup{Re}(s)>k+1$ by comparison with $D_{F,\mathcal{P}_{k,D,r}}(s)$ (see Lemma 6.1). Let also
@@
We then have that
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)D_{F,\xi_{i},\mathcal{P}}(s),$ (13)
where $D_{F,\xi_{i},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}A(N\xi_{i})N^{-s}$.
@@ 8. Relation to Sugano’s Theorem
In order to now obtain an Euler product, we will make use of the main theorem of Sugano in his paper [27]. We first need some setup.
@@
We remind ourselves here that $G_{\infty}^{0}$ acts transitively on $\mathcal{H}_{S}$ (cf. Section 2). Let $\mathcal{Z}_{0}$ denote any point of $\mathcal{H}_{S}$ with real part $0$ and denote by $K_{\infty}$ its stabiliser in $G_{\infty}^{0}$. Then, we have that $G_{\infty}^{0}/K_{\infty}\cong\mathcal{H}_{S}$.
@@
$\bm{F}(g_{f};\mathcal{Z}):=\bm{F}(g_{\infty}g_{f})j(g_{\infty},\mathcal{Z}_{0})^{k},$
@@
$\bm{F}(g_{f};\gamma\langle\mathcal{Z}\rangle)=j(\gamma,\mathcal{Z})^{k}\bm{F}(g_{f};\mathcal{Z})$
for all $\gamma\in\Gamma(g_{f})$ and $\mathcal{Z}\in\mathcal{H}_{S}$. Now, if $X\in V_{0}$, define the element $\gamma_{X}\in G_{\mathbb{Q}}$ by
$\gamma_{X}:=\begin{pmatrix}1&-X^{t}S_{0}&-\frac{1}{2}S_{0}[X]\\ 0&1_{n+2}&X\\ 0&0&1\end{pmatrix}$.
Now, the holomorphic function $\bm{F}(g_{f};\mathcal{Z})$ is invariant under $\mathcal{Z}\longmapsto\mathcal{Z}+X$ for $X$ in the lattice
@@
Hence, every such function admits a Fourier expansion of the form
$\bm{F}(g_{f};\mathcal{Z})=\sum_{\begin{subarray}{c}r\in\hat{L}(g_{f})\\ ir\in\mathcal{H}_{S}\end{subarray}}a(g_{f};r)e(2\phi_{0}(r,Z)),$
@@
Finally, let us introduce adelic Fourier coefficients. Let $\chi=\prod_{p\leq\infty}\chi_{p}$ be a character of $\mathbb{Q}_{\mathbb{A}}$ such that $\chi_{\mid\mathbb{Q}}=1$ and $\chi_{\infty}(x)=e(x)$ for all $x\in\mathbb{R}$. For $\eta\in V_{0}$ and $g\in G_{\mathbb{A}}$, we define
$\bm{F}_{\chi}(g;\eta):=\int_{V_{0}\backslash V_{\mathbb{A}}}\bm{F}(\gamma_{\chi}g)\chi\left(-2\phi_{0}(\eta,X)\right)dX.$
@@
$\bm{F}_{\chi}(g_{\infty}g_{f};\eta)=a(g_{f};\eta)j(g_{\infty};\mathcal{Z}_{0})^{-k}e(2\phi_{0}(\eta,g_{\infty}\langle\mathcal{Z}_{0}\rangle)).$ (14)
We then have from the above definitions (see [27, (1.14)]):
@@
We now want to show that there is a bijection between the spaces $\mathfrak{S}(K_{f})$ and $S_{k}(\Gamma_{S})$. This is true because of the following Lemma.
Lemma 8.1.
We have $G_{\mathbb{A}}=G_{\mathbb{Q}}G_{\infty}^{0}K_{f}$.
Proof.
From the proof of [27, Lemma 1], we have $G_{\mathbb{A}}=G_{\mathbb{Q}}G_{\mathbb{A},f}^{*}G_{\infty}^{0}K_{f}$, where we view $G^{*}$ as a subgroup of $G$ via $g\longmapsto\textup{diag}(1,g,1)$. But, from Lemma 7.2, we have $G_{\mathbb{A}}^{*}=G_{\mathbb{Q}}^{*}G_{\infty}^{*}K_{f}^{*}$, where we now define
$K_{f}^{*}:=\prod_{p<\infty}G^{*}(\mathbb{Z}_{p}).$
From this, we obtain $G_{\mathbb{A},f}^{*}=G_{\mathbb{Q}}^{*}K_{f}^{*}$. Therefore, because elements of $K_{f}^{*}$ and $G_{\infty}^{0}$ commute, we obtain
$G_{\mathbb{A}}=G_{\mathbb{Q}}G_{\mathbb{Q}}^{*}G_{\infty}^{0}K_{f}^{*}K_{f}=G_{\mathbb{Q}}G_{\infty}^{0}K_{f},$
as required. ∎
The bijection is now given by $\bm{F}\longmapsto F(\mathcal{Z}):=\bm{F}(g_{\infty})j(g_{\infty},\mathcal{Z}_{0})^{k}$, where $g_{\infty}\in G_{\infty}^{0}$ is chosen so that $g_{\infty}\langle\mathcal{Z}_{0}\rangle=\mathcal{Z}$.
Let now $g_{f}=(id,id,\cdots)$, which we denote by just $id$. It then follows that $F(\mathcal{Z})=\bm{F}(id,\mathcal{Z}),\,\,\forall\mathcal{Z}\in\mathcal{H}_{S}$. In that case, we have $L(g_{f})=L_{0}=\mathbb{Z}^{n+2}$. Now, $a(id,r)=A(r)$ for all $r\in L_{0}^{*}$. Hence
@@
for all $i$, where $\xi_{i}$’s are the representatives for $L_{0}\left[-\frac{D}{q},\frac{1}{2}\mathbb{Z}\right]/\Gamma(L_{0})$, as in (12).
Fix now a complete system of representatives $\{u_{i}\}_{i=1}^{h}$ for $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$, corresponding to the $\xi_{i}^{\prime}s$ of (12), in the sense of Proposition 7.2. Assume also that $u_{i,\infty}=1$ for all $i=1,\cdots,h$. We then prove the following Lemma.
@@
Let $\xi$ be as we have specified it right after Lemma 7.2. If now $\xi_{i}$’s and $u_{i}$’s are as above, we have
@@
By the definition of the $\xi_{i}$’s and Witt’s theorem ([24, Lemma 1.5 (ii)]), there is $\alpha\in G_{\mathbb{Q}}^{*}$ such that $\xi_{i}=\alpha\xi$. By the correspondence given by Shimura in Proposition 7.2, we get $\alpha^{-1}L_{0}=u_{i}L_{0}$. This then implies $\alpha u_{i}L_{0}=L_{0}$, so $\alpha u_{i}\in C$ (we remind ourselves that $C=\{x\in G^{*}_{\mathbb{A}}\mid xL_{0}=L_{0}\}$). By definition, we then get $\alpha u_{i}\in K_{f}$. Hence, we obtain
$\bm{F}_{\chi}(u_{i};\xi)=\bm{F}_{\chi}(u_{i};\alpha^{-1}\xi_{i})=\bm{F}_{\chi}\left(\textup{diag}(1,\alpha,1)u_{i};\xi_{i}\right)=\bm{F}_{\chi}\left(\left(\textup{diag}(1,\alpha,1),id\right);\xi_{i}\right),$
where $\left(\textup{diag}(1,\alpha,1),id\right)$ denotes the element of $G_{\mathbb{A}}$ with infinity part $\textup{diag}(1,\alpha,1)$. The second equality above follows from the second bullet point and the third equality from the first bullet point, just before Lemma 8.1. By the property of $\bm{F}_{\chi}$ in (14), we obtain
$a(u_{i};\xi)e\left(2\phi_{0}(\xi,\mathcal{Z}_{0})\right)=a(id,\xi_{i})j\left(\textup{diag}(1,\alpha,1),\mathcal{Z}_{0}\right)^{-k}e\left(2\phi_{0}\left(\xi_{i},\textup{diag}(1,\alpha,1)\langle\mathcal{Z}_{0}\rangle\right)\right).$
But $j\left(\textup{diag}(1,\alpha,1),\mathcal{Z}_{0}\right)=1$ and $\textup{diag}(1,\alpha,1)\langle\mathcal{Z}_{0}\rangle=\alpha\mathcal{Z}_{0}$ which then gives that the right hand side above equals
$a(id;\xi)e\left(2\phi_{0}(\xi_{i},\alpha\mathcal{Z}_{0})\right)=a(id;\xi)e\left(2\phi_{0}(\alpha\xi,\alpha\mathcal{Z}_{0})\right)=a(id;\xi)e\left(2\phi_{0}(\xi,\mathcal{Z}_{0})\right),$
@@
For each prime $p$ and $g_{f}\in G_{\mathbb{A},f}$, let
$M(g_{f};\xi)_{p}:=H(\xi)_{p}\cap g_{f}K_{f}g_{f}^{-1}\textup{ and }M(g_{f};\xi)_{f}:=\prod_{p}M(g_{f};\xi)_{p}.$
Define then $e(\xi)_{i}:=\#\{H(\xi)_{\mathbb{Q}}\cap M(u_{i}g_{f};\xi)_{f}\}$ for $1\leq i\leq h$ and $\displaystyle{\mu(\xi):=\sum_{i=1}^{h}e(\xi)^{-1}_{i}}$.
Let also $V(g_{f};\xi)$ the space of $\mathbb{C}$-valued functions on $H(\xi)_{\mathbb{A}}$, which are left $H(\xi)_{\mathbb{Q}}$ and right $H(\xi)_{\infty}M(g_{f};\xi)_{f}$ invariant. We now have the following Theorem, which follows from [27, Theorem 1].
@@
Assume $F\in S_{k}(\Gamma_{S})$ corresponds to $\bm{F}\in\mathfrak{S}_{k}(K_{f})$, as above. Assume $\bm{F}$ is a simultaneous eigenfunction of the Hecke pairs $\mathcal{H}_{p}=(G_{p},K_{p})$ for all $p$ (see [27, Section 2]). We also assume $A(\xi)\neq 0$ and $g_{f}=id$. Then, there is a finite set of primes $\mathcal{P}$, such that if $f\in V(g_{f};\xi)$ is a simultaneous eigenfunction of the Hecke algebras $\mathcal{H}_{p}^{\prime}:=(H(\xi)_{p},M(g_{f};\xi)_{p})$ for all $p\notin\mathcal{P}$, we have
@@
$\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}\mu(\xi)^{-1}\sum_{i=1}^{h}A(N\xi_{i})\frac{\overline{f}(u_{i})}{e(\xi)_{i}}N^{-(s+k-\frac{n+2}{2})}=\left(\mu(\xi)^{-1}\sum_{i=1}^{h}A(\xi_{i})\frac{\overline{f}(u_{i})}{e(\xi)_{i}}\right)\times\\ \times L_{\mathcal{P}}(F;s)L_{\mathcal{P}}\left(\overline{f};s+1/2\right)^{-1}\times(\textup{zeta})^{-1}(s),$
where $(\textup{zeta})(s):=\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s)&\textup{if }n\textup{ even}\end{cases}$.
@@
Here, $L(-,s)$ denotes the standard $L$-function attached to orthogonal modular forms, as this is defined in [27, Paragraph 4-1]. Also, for any $L$-function, we write $L_{\mathcal{P}}$ for the Euler product not containing the primes in $\mathcal{P}$.
Proof.
This follows by Sugano’s main Theorem in [27, Theorem 1]. In our setting, we take $g_{f}=(id,id,\cdots)$ and then substitute $a(u_{i};N\xi)$ with $A(N\xi_{i})$ in the original form of [27, Theorem 1], due to Lemma 8.2. We also note that in this case $H(\xi)_{\infty}M(g_{f};\xi)_{f}=H(\xi)_{\mathbb{A}}\cap C$, where $C$ is defined in (11). The set of primes $\mathcal{P}$ contains all the primes contained in the set $\mathcal{P}_{2}$ of [27, Theorem 1] and the finite set of primes $p$ for which $\partial_{p}\neq 0$ or $\partial^{\prime}_{p}\neq 0$, where $\partial_{p},\partial_{p}^{\prime}$ are defined in [27, Theorem 1]. We note that in our case, $L_{p}$ is maximal for all $p$, so $\mathcal{P}_{1}$ in [27, Theorem 1] is empty. ∎
From now on, we fix $g_{f}=id$. For any $f\in V(g_{f};\xi)$, we set
$\widetilde{f}(u_{i}):=\frac{\overline{f}(u_{i})}{e(\xi)_{i}}\mu(\xi)^{-1}\textup{, }i=1,\cdots,h\,\ \textup{and }\displaystyle{A_{f}:=\sum_{i=1}^{h}\widetilde{f}(u_{i})A(\xi_{i})}.$ (15)
The formula in Theorem 8.1 then becomes
@@
$(\textup{zeta})(s)\times L_{\mathcal{P}}\left(\overline{f};s+1/2\right)\sum_{i=1}^{h}\widetilde{f}(u_{i})D_{F,\xi_{i},\mathcal{P}}\left(s+k-(n+2)/2\right)=A_{f}L_{\mathcal{P}}(F;s),$
@@
where $D_{F,\xi_{i},\mathcal{P}}(s)$ is the Dirichlet series appearing in Proposition 7.3. This is true for any simultaneous eigenfunction $f$ (hence $\widetilde{f}$ as well) of the Hecke algebras $H_{p}^{\prime}=(H(\xi)_{p},M(g_{f};\xi)_{p})$ for all $p\notin\mathcal{P}$. Our aim is to invert it so that we solve for $D_{F,\xi_{i},\mathcal{P}}(s)$.
From the definition of the $u_{i}^{\prime}s$ just before Lemma 8.2, we have that
@@
where $D:=H(\xi)_{\infty}M(g_{f};\xi)_{f}$. We note $D$ is an open subgroup of $H(\xi)_{\mathbb{A}}$ and $D\cap H(\xi)_{f}$ is compact. By [24, Lemma 17.6], there is a correspondence